The Ramapo Company produces two products, Blinks and Dinks. They are manufactured in two departments, Fabrication and Assembly. Data for the products and departments are listed below.


Product
Number of
Units
Labor Hours
Per Unit
Machine Hours
Per Unit
Blinks 1,064 3 4
Dinks 2,300 6 8

All of the machine hours take place in the Fabrication department, which has an estimated overhead of $109,600. All of the labor hours take place in the Assembly department, which has an estimated total overhead of $106,800.

The Ramapo Company uses a single overhead rate to apply all overhead costs based on labor hours. What is the overhead cost per unit for Dinks?

a.$61.15

b.$27.95

c.$320.42

d.$76.44
Answer #1

Solution:

Estimated labor hours = (1064*3) + (2300*6) = 16992 hours

Predetermined overhead rate = Estimated overhead / Estimated direct labor hours

= ($109,600 + $106,800) / 16992 = $12.74 per DLH

Overhead cost per unit for Dinks = Labor hour per unit * Overhead rate per labor hour

= 6 * $12.74

= $76.44 per unit

Hence option d is correct.